If a 10-0 L sample of gas at STP is heated until its volume is 2`1 L and it’s temperature is 451 Celsius degrees, what is the new pressure in kPa
1
Expert's answer
2020-09-05T14:42:41-0400

T2 = 451 + 273.15 = 724.15 K

P1V1/T1 = P2V2/T2

"101.325 kPa * 10 L \/ 273.15 K = P2 * 2.1 L \/ 724.15 K"

P2 = 1279.2 kPa